# A Fine-Grained Attention and Geometric Correspondence Model for Musculoskeletal Risk Classification in Athletes Using Multimodal Visual and Skeletal Features

### v1

@@ Abstract
Musculoskeletal disorders pose significant risks to athletes, and assessing risk early is important for prevention. However, most existing methods are designed for controlled settings and fail to reliably assess risk in complex environments due to their reliance on a single type of data. This research proposes ViSK-GAT (Visual-Skeletal Geometric Attention Transformer), a novel multimodal deep learning framework designed to classify musculoskeletal risk using visual and skeletal coordinate-based features. In addition, a custom multimodal dataset is constructed by combining visual data and skeletal coordinates for risk assessment. Each sample is labeled into eight risk categories based on the Rapid Entire Body Assessment system. ViSK-GAT combines a Residual Block with a Lightweight Transformer Block to learn spatial and temporal dependencies jointly. It incorporates two novel modules: the Fine-Grained Attention Module (FGAM), which enables precise inter-modal feature refinement through cross-attention between visual and skeletal inputs, and the Multimodal Geometric Correspondence Module (MGCM), which enhances cross-modal coherence by aligning image features with coordinate-based representations. ViSK-GAT achieved strong performance with validation and test accuracies of 93.55% and 93.89%, respectively; a precision of 93.86%; an F1 score of 93.85%; and Cohen’s Kappa and Matthews Correlation Coefficient of 93%. The regression results also indicated a low Root Mean Square Error of the predicted probability distribution of 0.1205 and a corresponding Mean Absolute Error of 0.0156. Compared to nine popular transfer learning backbones, ViSK-GAT consistently outperformed previous methods. The ViSK-GAT model advances artificial intelligence implementation and application, transforming musculoskeletal risk classification and enabling impactful early interventions in sports.
Keywords: Deep Learning; Artificial Intelligence; Musculoskeletal Risk; Early Intervention; Multimodal Framework
@@ 1  Introduction
Musculoskeletal disorders (MSD) are among the main concerns in athletic performance, which pose significant risks that can affect movement efficiency, endurance, and long-term health [1]. These disorders arise from repetitive strain, improper biomechanics, and sustained postural imbalances, leading to muscle fatigue, joint stress, and chronic injuries [2]. Given these challenges, posture plays an important role in athlete’s performance, endurance, and injury prevention. Proper posture improves balance, stability, and coordination, key factors in executing precise movements. In contrast, poor posture can cause muscle imbalances and an increased risk of musculoskeletal injuries, such as strains, sprains, and joint stress [3]. Therefore, effective movement management is essential not only for athletes but for individuals of all ages and levels of activity, as it supports long-term health, safety, and overall well-being [4]. Researchers have continuously searched for ways to design better training environments and develop techniques that recognize movements [5]. However, addressing these risks requires a proactive approach to posture monitoring and correction, where sports ergonomics becomes essential. Despite their levels of athleticism, many athletes experience postural disturbances due to repetitive training [6, 7], eventually leading to musculoskeletal degeneration and chronic conditions [8, 9].
@@
Recently, posture assessment has increasingly transitioned into the automation phase. For example, sensor-based methods [18, 19, 20] provide the advantage of continuous monitoring, while computer vision-based approaches [21, 22, 23] offer non-invasive alternatives. More recently, multimodal methods [24, 25, 26] have emerged to integrate complementary data types for posture evaluation.
Although existing models have shown success in workplace ergonomics and human activity recognition, they often fail to address the unique challenges present in sports environments. Many systems struggle with real-time adaptability, occlusions during dynamic movements, and the effective integration of multimodal data sources. Furthermore, issues such as signal instability, calibration sensitivity, and discomfort caused by monitoring devices can interfere with the natural performance of an athlete [27, 28]. Capturing dynamic posture changes in fast-paced sports remains difficult, as environmental factors such as occlusion and limited field of view can hinder reliable evaluations [21, 22]. Although some approaches aim to combine multiple data sources, they often fail to effectively align and integrate various types of information, especially under rapidly changing conditions [24, 25]. Furthermore, many models were designed for controlled settings, which limits their applicability to the dynamic real-world environment [29, 26]. These challenges highlight the need for a more adaptable system capable of providing precise real-time posture assessments in athletic environments.
@@
Thus, to address these critical limitations, we propose ViSK-GAT (Visual-Skeletal Geometric Attention Transformer), a multimodal framework designed for robust and interpretable posture risk assessment in athletic environments. Unlike conventional approaches (see Figure 1) that struggle under real-world conditions, ViSK-GAT effectively combines high-level visual features with fine-grained skeletal data, enabling accurate posture classification even during dynamic movements and partial occlusion. This is achieved through two key innovations: the Fine-Grained Attention Module (FGAM), which enhances cross-modal feature interaction, and the Multimodal Geometric Correspondence Module (MGCM), which ensures precise alignment between heterogeneous data sources using attention-based geometric reasoning. Our extensive evaluations highlight the superiority of ViSK-GAT, which achieves a test accuracy of 93.89%, significantly outperforming all tested backbone variants, including DenseNet121 (88.06%), ResNet101 (88.39%), and EfficientNet-B0 (87.42%), even when these models are paired with identical fusion and classification pipelines. Furthermore, ViSK-GAT attains the lowest validation and test losses (0.4376 and 0.4808, respectively), indicating improved generalization and model stability. Compared to previous studies on ergonomic risk assessment, ViSK-GAT demonstrates leading and competitive performance. Although prior models using ergonomic metrics such as OWAS and RULA report test accuracies ranging from 76.10% to 93.00%, ViSK-GAT surpasses these with a test accuracy of 93.89%, also using the more fine-grained and anatomically detailed REBA scoring system, which is better suited for dynamic posture analysis in sports. These results collectively establish ViSK-GAT as a great solution, both in terms of accuracy and architectural innovation, for multimodal ergonomic risk assessment, particularly in the complex and fast-paced conditions of athletic environments.
@@ 2.1  Vision-based methods
Vision-based methods mostly use deep learning networks to analyze human activities with potential MSD risks in a non-invasive way. In the study of Li et al. [21], a vision-based real-time method was proposed to assess postural risk factors for musculoskeletal disorders using CNN feature extraction to generate 2D postures. The algorithm achieved 93% accuracy in detecting RULA action levels. On the other hand, Kim et al. [22] adopted an open-source technology named OpenPose to develop a human pose estimation system. The system counted the angles of each part of the body from a two-dimensional perspective to assess the level of risk of posture. Seo et al. [23] introduced a computer vision-based method for the automatic identification of construction worker postures to assess ergonomic risk. Using classification techniques to learn from various postures depicted in virtual images, they achieved 89% precision in classifying diverse postures in the images. In another study, Dzeng et al. [30] proposed an approach that automatically tracks and categorizes postures based on OWAS using skeleton-based motion data to analyze ergonomic risks. The posture identification accuracies exceeded 90% for most activities, with up to 88.5% in the OWAS assessment. On the same assessment (i.e, OWAS), Yan et al. [31] achieved a maximum accuracy of 89.2% using three posture classifiers for construction hazard prevention. Many researchers have also proposed ergonomic assessment based on skeletons extracted either from 2D images from an ordinary camera or 3D images produced by RGB-D sensors [32, 33, 31, 34].
@@ 2.2  Multimodal-based methods
For instance, Xiahou et al. [24] proposed a feature-level fusion framework that integrates signals from various sensors using deep learning models such as multilayer perceptrons (MLPs), recurrent neural networks (RNNs), and long short-term memory (LSTM) networks. Their LSTM-based model maintained over 90% accuracy even under adverse environmental conditions. Similarly, Zhao et al. [25] designed a multimodal bioinformation fusion algorithm that extracted motion features from both hand and leg movements using wavelet packet decomposition and sample entropy. Their method achieved recognition rates consistently above 95%. Several studies have also combined visual and sensor data for posture and behavior recognition. For instance, Zhang et al. [26] introduced a deep learning framework that fuses infrared and pressure map data to recognize sitting postures. Using modality-specific backbones and a cross-modal self-attention mechanism, they achieved an F1-score of 93.08%. Furthermore, Palash et al. [35] introduced a modular two-stream CNN and RNN framework that fuses facial expressions, body posture, and gait features early, while leveraging situational knowledge to enhance multimodal emotion recognition. Meanwhile, Qi et al. [36] introduced a federated learning-based multimodal framework for fall detection, where time-series data from wearable sensors and visual data from cameras are fused at the input level using a Gramian Angular Field transformation. Their model reached up to accuracies of 99.927% and 89.769% for binary and multi-class fall activity recognition tasks, respectively. Similarly, Liu et al. [37] developed a hydrogel-based system for monitoring respiration and recognizing finger-pressing postures. Even though they achieved a posture recognition accuracy of 99.259%, their work primarily relied on a wearable sensor system, which often adds additional inconvenience to the users.
@@ 3.1  MusDis-Sports dataset
Our MusDis-Sports dataset comprises 8,989 annotated samples, each consisting of an image of a human subject in a sports-related posture, the corresponding 2D skeletal keypoints, and an ergonomically informed class label representing the REBA-based risk category. Figure 2 visualizes some images from the constructed dataset. The dataset can be accessed through the footnoted GitHub repository [^1] https://github.com/mak-raiaan/MusDis-Sports_Dataset.
Figure 2: Samples from the proposed dataset.
@@ 3.1.1  Image and skeleton coordinate construction
$\theta\_{ABC}=\cos^{-1}\left(\frac{(\vec{BA}\cdot\vec{BC})}{|\vec{BA}|\cdot|\vec{BC}|}\right)$ (2)
@@
1:Input: images $\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$ each image $I\in\mathcal{I}$
2:Apply MediaPipe to keypoints: $\mathcal{K}=\{(x_{i},y_{i},v_{i})\}_{i=1}^{33}$
3:Initialize empty keypoint list K{K} each keypoint $(x_{i},y_{i},v_{i})\in\mathcal{K}$ $v_{i}\geq 0.5$
4:Rescale $(x_{i},y_{i})$ to original image size
5:Append to K{K} each joint triplet $(A,B,C)$ in K{K}
6:vectors: $\vec{BA}\leftarrow A-B$, $\vec{BC}\leftarrow C-B$
7:angles: compute angle $\theta_{ABC}$ each body segment
8:v_angel: vertical inclination angle $\theta$
9:Overlay keypoints and skeletal edges on image II
10:Save the annotated image
@@ 3.1.2  REBA labeling
We computed ergonomic risk scores using a structured implementation of the Rapid Entire Body Assessment (REBA) protocol, following the methodology proposed by Hignett and McAtamney [10]. It is a standardized tool for evaluating the risk of musculoskeletal disorders based on body posture and has been widely used in ergonomics and occupational health studies [43]. The process involves segmenting the human body into functional regions (e.g., trunk, neck, legs, upper arms, lower arms, and wrists) and assigning each part a discrete risk score based on the estimated joint angles.
@@
REBA Class Samples
Class 1 1083
Class 2 1213
Class 3 1181
Class 4 1332
Class 5 1144
Class 6 1117
Class 7 888
Class 8 1031
Total 8989
Our study specifically focuses on the 6 body parts: neck, trunk, wrist, upper arm, lower arm, and legs. As our dataset consists of single-frame images without dynamic tasks or load handling, we fixed both the load and activity modifiers to zero. The final Group C value was treated as the REBA scores; and based on the final calculated scores, we defined 8 classes (see Table 1) where classes with a higher number (i.e., 4+) indicate riskier postures. The REBA score assignment process is briefed in Algorithm 2.
@@
[…]
2:use REBA lookup to compute $g_{A}$
[…]
4:use REBA lookup to compute $g_{B}$
@@
[…]
12:Output: final REBA score, $S_{\text{REBA}}$

### v2

@@ Abstract
Musculoskeletal disorders pose significant risks to athletes, and assessing risk early is important for prevention. However, most existing methods are designed for controlled settings and fail to reliably assess risk in complex environments due to their reliance on a single type of data. This research introduces ViSK-GAT (Visual-Skeletal Geometric Attention Transformer), a novel multimodal deep learning framework that classifies musculoskeletal risk using both visual and skeletal coordinate-based features. A custom multimodal dataset (MusDis-Sports) was created by combining images and skeletal coordinates, with each sample labeled into eight risk categories based on the Rapid Entire Body Assessment (REBA) system. ViSK-GAT integrates two innovative modules: the Fine-Grained Attention Module (FGAM), which refines inter-modal features via cross-attention between visual and skeletal inputs, and the Multimodal Geometric Correspondence Module (MGCM), which enhances cross-modal alignment between image features and coordinates. The model achieved robust performance, with all key metrics exceeding 93%. Regression results also indicated a low RMSE of 0.1205 and MAE of 0.0156. ViSK-GAT consistently outperformed nine popular transfer learning backbones and showed its potential to advance AI-driven musculoskeletal risk assessment and enable early, impactful interventions in sports.
Keywords: Deep Learning; REBA; Musculoskeletal Risk; Posture Recognition; Multimodal Framework; Fine-Grained Attention
@@ 1  Introduction
Musculoskeletal disorders (MSD) are among the main concerns in athletic performance, which pose significant risks that can affect movement efficiency, endurance, and long-term health [1]. These disorders arise from repetitive strain, improper biomechanics, and sustained postural imbalances, leading to muscle fatigue, joint stress, and chronic injuries [2]. Given these challenges, posture plays an important role in athletes’ performance, endurance, and injury prevention. Proper posture improves balance, stability, and coordination, key factors in executing precise movements. In contrast, poor posture can cause muscle imbalances and an increased risk of musculoskeletal injuries, such as strains, sprains, and joint stress [3]. Therefore, effective movement management is essential not only for athletes but for individuals of all ages and levels of activity, as it supports long-term health, safety, and overall well-being [4]. Researchers have continuously sought ways to design more effective training environments and develop techniques that recognize movements [5]. However, addressing these risks requires a proactive approach to posture monitoring and correction, where sports ergonomics becomes essential. Despite their levels of athleticism, many athletes experience postural disturbances due to repetitive training [6, 7], eventually leading to musculoskeletal degeneration and chronic conditions [8, 9].
@@
Recently, posture assessment has increasingly transitioned into the automation phase. For example, sensor-based methods [18, 19, 20] provide the advantage of continuous monitoring, while computer vision-based approaches [21, 22, 23, 45] offer non-invasive alternatives. More recently, multimodal methods [25, 26, 27] have emerged to integrate complementary data types for posture evaluation. Although existing models have shown success in workplace ergonomics and human activity recognition, they often fail to address the unique challenges present in sports environments. Many systems struggle with real-time adaptability, occlusions during dynamic movements, and the effective integration of multimodal data sources. Furthermore, issues such as signal instability, calibration sensitivity, and discomfort caused by monitoring devices can interfere with the natural performance of an athlete [28, 29]. Capturing dynamic posture changes in fast-paced sports remains difficult, as environmental factors such as occlusion and limited field of view can hinder reliable evaluations [21, 22]. Although some approaches aim to combine multiple data sources, they often fail to effectively align and integrate various types of information, especially under rapidly changing conditions [25, 26]. Furthermore, many models were designed for controlled settings, which limits their applicability to the dynamic real-world environment [30, 27]. These challenges highlight the need for a more adaptable system capable of providing precise real-time posture assessments in athletic environments.
@@
Thus, to address these critical limitations, we propose ViSK-GAT (Visual-Skeletal Geometric Attention Transformer), a multimodal framework designed for robust and interpretable posture risk assessment in athletic environments. While cross-attention is an established mechanism for multimodal fusion, its standard application often treats modalities as monolithic blocks for global alignment. This can be suboptimal for fine-grained postural analysis, where specific local image features (e.g., a bent elbow) must be precisely correlated with specific skeletal keypoints. To address this, ViSK-GAT introduces a structured, two-stage attention pipeline. The Fine-Grained Attention Module (FGAM) first performs intra-modal refinement, enhancing discriminative local features within the image data before fusion. This prepared representation is then processed by the Multimodal Geometric Correspondence Module (MGCM), which performs a spatially-aware cross-modal alignment. Unlike generic cross-attention, MGCM is designed explicitly to bridge the domain gap between visual texture and geometric coordinate spaces, ensuring that the model’s fusion is grounded in the biomechanical reality of the human body.
Furthermore, unlike conventional approaches (see Figure 1) that struggle under real-world conditions, ViSK-GAT effectively combines high-level visual features with fine-grained skeletal data, enabling accurate posture classification even during dynamic movements and partial occlusion. Our extensive evaluations highlight the superiority of ViSK-GAT, which achieves a test accuracy of 93.89%, significantly outperforming state-of-the-art backbones, even when these models are paired with identical fusion and classification pipelines. Compared to previous studies on ergonomic risk assessment, ViSK-GAT demonstrates leading and competitive performance. Although prior models using ergonomic metrics such as OWAS and RULA report test accuracies ranging from 76.10% to 93.00%, ViSK-GAT surpasses these with a test accuracy of 93.89%, also using the more fine-grained and anatomically detailed REBA scoring system, which is better suited for dynamic posture analysis in sports. These results collectively establish ViSK-GAT as a great solution, both in terms of accuracy and architectural innovation, for multimodal ergonomic risk assessment, particularly in the complex and fast-paced conditions of athletic environments.
@@
Figure 2: Sample images from the constructed MusDis-Sports dataset showing athletes in various sports-related postures with overlaid 2D skeletal keypoints extracted using the MediaPipe framework. These samples are used for training and evaluating musculoskeletal risk classification.
@@ 2.1  Vision-based methods
Vision-based methods mostly use deep learning networks to analyze human activities with potential MSD risks in a non-invasive way. In the study of Li et al. [21], a vision-based real-time method was proposed to assess postural risk factors for musculoskeletal disorders using CNN feature extraction to generate 2D postures. The algorithm achieved 93% accuracy in detecting RULA action levels. On the other hand, Kim et al. [22] adopted an open-source technology named OpenPose to develop a human pose estimation system. The system counted the angles of each part of the body from a two-dimensional perspective to assess the level of risk of posture. Seo et al. [23] introduced a computer vision-based method for the automatic identification of construction worker postures to assess ergonomic risk. Using classification techniques to learn from various postures depicted in virtual images, they achieved 89% precision in classifying diverse postures in the images. In another study, Dzeng et al. [31] proposed an approach that automatically tracks and categorizes postures based on OWAS using skeleton-based motion data to analyze ergonomic risks. The posture identification accuracies exceeded 90% for most activities, with up to 88.5% in the OWAS assessment. On the same assessment (i.e, OWAS), Yan et al. [32] achieved a maximum accuracy of 89.2% using three posture classifiers for construction hazard prevention. Many researchers have also proposed ergonomic assessment based on skeletons extracted either from 2D images from an ordinary camera or 3D images produced by RGB-D sensors [32, 33, 34, 35, 45].
@@ 2.2  Multimodal-based methods
For instance, Xiahou et al. [25] proposed a feature-level fusion framework that integrates signals from various sensors using deep learning models such as multilayer perceptrons (MLPs), recurrent neural networks (RNNs), and long short-term memory (LSTM) networks. Their LSTM-based model maintained over 90% accuracy even under adverse environmental conditions. Similarly, Zhao et al. [26] designed a multimodal bioinformation fusion algorithm that extracted motion features from both hand and leg movements using wavelet packet decomposition and sample entropy. Their method achieved recognition rates consistently above 95%. Several studies have also combined visual and sensor data for posture and behavior recognition. For instance, Zhang et al. [27] introduced a deep learning framework that fuses infrared and pressure map data to recognize sitting postures. Using modality-specific backbones and a cross-modal self-attention mechanism, they achieved an F1-score of 93.08%. Furthermore, Palash et al. [36] introduced a modular two-stream CNN and RNN framework that fuses facial expressions, body posture, and gait features early, while utilizing situational knowledge to enhance multimodal emotion recognition. Meanwhile, Qi et al. [37] introduced a federated learning-based multimodal framework for fall detection, where time-series data from wearable sensors and visual data from cameras are fused at the input level using a Gramian Angular Field transformation. Their model reached up to accuracies of 99.927% and 89.769% for binary and multi-class fall activity recognition tasks, respectively. Similarly, Liu et al. [38] developed a hydrogel-based system for monitoring respiration and recognizing finger-pressing postures. Even though they achieved a posture recognition accuracy of 99.259%, their work primarily relied on a wearable sensor system, which often adds additional inconvenience to the users.
@@ 3.1  MusDis-Sports dataset
Our MusDis-Sports dataset comprises 8989 annotated samples, each consisting of an image of a human subject in a sports-related posture, the corresponding 2D skeletal keypoints, and an ergonomically informed class label representing the REBA-based risk category. Figure 2 visualizes some images from the constructed dataset. The dataset can be accessed through the footnoted GitHub repository [^1] https://github.com/mak-raiaan/MusDis-Sports_Dataset.
@@ 3.1.1  Image and skeleton coordinate construction
$\theta_{ABC}=\cos^{-1}\left(\frac{(\vec{BA}\cdot\vec{BC})}{|\vec{BA}|\cdot|\vec{BC}|}\right)$ (2)
@@
1:Input: images $\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$
2:for each image $I\in\mathcal{I}$ do
3:  Apply MediaPipe to keypoints: $\mathcal{K}=\{(x_{i},y_{i},v_{i})\}_{i=1}^{33}$
4:  Initialize empty keypoint list $\mathbf{K}$
5:  for each keypoint $(x_{i},y_{i},v_{i})\in\mathcal{K}$ do
6:   if $v_{i}\geq 0.5$ then
7:     Rescale $(x_{i},y_{i})$ to original image size
8:     Append to $\mathbf{K}$
9:   end if
10:  end for
11:  for each joint triplet $(A,B,C)$ in $\mathbf{K}$ do
12:   vectors: $\vec{BA}\leftarrow A-B$, $\vec{BC}\leftarrow C-B$
13:   angles: compute angle $\theta_{ABC}$
14:  end for
15:  for each body segment do
16:   v_angel: vertical inclination angle $\theta$
17:  end for
18:  Overlay keypoints and skeletal edges on image II
19:  Save the annotated image
20:end for
@@ 3.1.2  REBA labeling
We computed ergonomic risk scores using a structured implementation of the REBA protocol, following the methodology proposed by Hignett and McAtamney [10]. It is a standardized tool for evaluating the risk of musculoskeletal disorders based on body posture and has been widely used in ergonomics and occupational health studies [44]. The process involves segmenting the human body into functional regions (e.g., trunk, neck, legs, upper arms, lower arms, and wrists) and assigning each part a discrete risk score based on the estimated joint angles.
@@
Our study specifically focuses on the 6 body parts: neck, trunk, wrist, upper arm, lower arm, and legs. As our dataset consists of single-frame images without dynamic tasks or load handling, we fixed both the load and activity modifiers to zero. The final Group C value was treated as the REBA scores, and based on the final calculated scores, we defined 8 classes, where classes with a higher number (i.e., 4+) indicate riskier postures. Table 1 reports the dataset samples and ratio in terms of eight classes.
@@
REBA Class Samples Ratio (approx.)
Class 1 1083 12.05%
Class 2 1213 13.49%
Class 3 1181 13.14%
Class 4 1332 14.82%
Class 5 1144 12.73%
Class 6 1117 12.43%
Class 7 888 09.88%
Class 8 1031 11.47%
Total 8989 100%
Although some REBA protocol relies on 3D angular measurements, we employed a 2D approximation for broader applicability in conventional available image data that preserves the relative ranking of joint stresses. The resulting classes, therefore, represent postural intensity levels derived from 2D angular displacements. This reinterpretation ensures the labels remain meaningful even within monocular visual data. To ensure labeling reliability despite the automatic nature of REBA score generation, (i) we used MediaPipe’s built-in visibility confidence scores to exclude low-certainty landmarks (vi < 0.7); (ii) for each body-region score, we analyzed internal consistency by correlating the derived angular magnitudes across overlapping joints (e.g., shoulder–elbow–wrist chains) and retained samples with consistent geometry; and (iii) Furthermore, to assess labeling stability, we performed a manual sanity check on a random subset of 500 images to visually confirm that the overlay of skeletal keypoints was accurate and that the assigned REBA class was plausible given the posture. We also processed the samples under slightly varied viewing scales and compared them using the intraclass correlation coefficient (ICC = 0.84). Our findings indicate that the automatically derived REBA labels are self-consistent within a reasonable tolerance for 2D estimation. The REBA score assignment process is briefed in Algorithm 2.
@@
1:Input: Scores: $s_{\text{neck}}$, $s_{\text{trunk}}$, $s_{\text{leg}}$, $s_{\text{upper}}$, $s_{\text{lower}}$, $s_{\text{wrist}}$
2:function GroupA($s_{\text{neck}},s_{\text{trunk}},s_{\text{leg}}$)
3:  use REBA lookup to compute $g_{A}$
4:  return $g_{A}$
5:end function
6:function GroupB($s_{\text{upper}},s_{\text{lower}},s_{\text{wrist}}$)
7:  use REBA lookup to compute $g_{B}$
8:  return $g_{B}$
9:end function
10:function GroupC($g_{A},g_{B}$)
11:  $S_{\text{REBA}}\leftarrow\text{table}[g_{A}][g_{B}]$
12:  return $S_{\text{REBA}}$
13:end function
@@
17:return $S_{\text{REBA}}$
18:Output: final REBA score, $S_{\text{REBA}}$